OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (anonymous):

2x+6-4=20 and 2x+6-4=-20 right?

OpenStudy (anonymous):

Wait no. Put 4 on right side?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

first move everything from left to right side except the absolute bar

OpenStudy (anonymous):

okay, so how do I do that?

Nnesha (nnesha):

do what ?

Nnesha (nnesha):

do you mean how to move -4 to the right side ?

OpenStudy (anonymous):

Move everything.

OpenStudy (anonymous):

Oh, so we just move -4?

Nnesha (nnesha):

there is only -4 how would you cancel out -4 from left side ?

OpenStudy (anonymous):

You can't.

Nnesha (nnesha):

yes but which action would take subtract/add/multiply or divide both sides by -4 ?

Nnesha (nnesha):

you*

OpenStudy (anonymous):

So 20-4

OpenStudy (anonymous):

Subtract?

Nnesha (nnesha):

hmm no it's already negative at left side so do opposite of subtraction

OpenStudy (anonymous):

Oh, addition. Okay.

Nnesha (nnesha):

right! add 4 both sides then you can separate that into 2 equations

OpenStudy (anonymous):

\[2x+6-4+4=20+4?\]

Nnesha (nnesha):

don't forget the absolute bar \[\left| 2x+6 \right| -4+4=20 +4\]

OpenStudy (anonymous):

Oh, okay. So then I solve right?

OpenStudy (anonymous):

2x+6=24

Nnesha (nnesha):

and 2nd equation ?

OpenStudy (anonymous):

2x+6-4+4=-20+4?

Nnesha (nnesha):

no first to get rid of everything except the absolute bar you got \[\left| 2x+6 \right|=24\] now separate the absolute bar into two equation

OpenStudy (anonymous):

How?

Nnesha (nnesha):

just like we did on the previous post \[\rm \left| a+b \right| = c ~~~~~~~\left| a+b\right| =- c \]

OpenStudy (anonymous):

2x+6=-20

Nnesha (nnesha):

how did you get -20?

OpenStudy (anonymous):

Woops.

Nnesha (nnesha):

remember ) first we need absolute bar itself on the left side 2nd) then we can spllit absolute into the equation with each sign (plus/negative)

OpenStudy (anonymous):

Isn't it -20?

Nnesha (nnesha):

no. what did you get when u added 4 both sides ?

OpenStudy (anonymous):

24

Nnesha (nnesha):

right so it should be 24 not 20

OpenStudy (anonymous):

so, 2x+6=-24?

Nnesha (nnesha):

right

OpenStudy (anonymous):

2x=-30

OpenStudy (anonymous):

x=-15

Nnesha (nnesha):

good

OpenStudy (anonymous):

So, 9 and -15?

Nnesha (nnesha):

did you get 9 for first equation ?

OpenStudy (anonymous):

Yes.

Nnesha (nnesha):

ok then. right!

OpenStudy (anonymous):

Thanks! :)

Nnesha (nnesha):

np